OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 2x+y+8 and passes through the point (-2,1)

OpenStudy (anonymous):

Me thinks you got typo

OpenStudy (anonymous):

Its on the sheet my teacher gave me

OpenStudy (anonymous):

oh wait i do its 2x+y=8 .

OpenStudy (anonymous):

Alright, so can you put that into slope intercept form?

OpenStudy (anonymous):

is that y=mx+b?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so then would it be y=2x+8?

OpenStudy (anonymous):

-2x y=-2x+8

OpenStudy (anonymous):

how do you show your work for this question?

OpenStudy (anonymous):

Just show the steps you took... from here you can see the slope of the line is -2, right?

OpenStudy (anonymous):

2x+y=8 -2x -2x y=-2x+8 thats it so far

OpenStudy (anonymous):

okay whats next?

OpenStudy (anonymous):

Use point slope form y-y1=m(x-x1) (x1,y1) is the point you have m=the slope which has to be the same as the other line since theyre parallel, -2 plug those in and simplify

OpenStudy (anonymous):

so you write it like this: y-1=m(x-(-2)?

OpenStudy (anonymous):

Forgot m y-1=-2(x-(-2))

OpenStudy (anonymous):

Now use distributive property on the right side, then add 1 to both sides

OpenStudy (anonymous):

so for the distrubuted side it would be 2x-4?

OpenStudy (anonymous):

the 2 in front is negative, y-1=-2x-4

OpenStudy (anonymous):

okay then you add 1 to both sides?

OpenStudy (anonymous):

Right, what's your final answer?

OpenStudy (anonymous):

y=2x-3. is that right?

OpenStudy (anonymous):

The negative keeps vanishing.... y=-2x-3

OpenStudy (anonymous):

so then that would be the final answer ^?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

thank you sooo much. can you help me with another one?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

ill supervise

OpenStudy (anonymous):

what is the slope of the line that passes through the points (3,-7) and (-5,-1)?

OpenStudy (anonymous):

Well, slope is \[\Delta y/\Delta x\]

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

So your points are (x1,y1) and (x2,y2) to find the slope you do (y2-y1)/(x2-x1)

OpenStudy (anonymous):

rise over run

OpenStudy (anonymous):

(-1-(-7))/(-5-3)=?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

uhm.... How do you find the answer?

OpenStudy (anonymous):

simplify the problem i just typed.. (-1+7)/(-5-3)

OpenStudy (anonymous):

how? I dont get it

OpenStudy (anonymous):

whats -1+7?

OpenStudy (anonymous):

do your pemdas

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yeah and whats -5-3?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

so, slope is 6/-8 which simplifies to?

OpenStudy (anonymous):

-3/4?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thank you so much(:

OpenStudy (anonymous):

You're welcome